OpenStudy (anonymous):

If the product of two consecutive odd integers is decreased by one third the lesser integer, the result is 250. Find the integers.

OpenStudy (agent47):

x*(x+2)-(x/3)=50

mathslover (mathslover):

let the first integer be x-1 then x+1and then they are decreased by (x-1)/3 therefore (x-1)(x+1)-(x-1)/3=250 x^2-1-(x-1)/3=250 (3x^2-3-x+1)=750 3x^2-x-2=750 can u solve it now

OpenStudy (anonymous):

Let the first odd integer be x And the other will be (x + 2).. x(x + 2) - x/3 = 250 \[x^2 + 2x - x/3 = 250\] \[3x^2 + 6x - x = 250*3\] \[3x^2 + 5x = 750\] \[3x^2 + 5x - 750 = 0\] Solve it by using Quadratic Formula...

OpenStudy (anonymous):

two very nice ways of solving the problem... :)

OpenStudy (agent47):

*had a typo lol i meant 250 not 50

OpenStudy (anonymous):

now there are three nice ways...:)

OpenStudy (anonymous):

Solve the following for n:\[(2n+1)(2n+3)-\frac{1}{3}(2n+1)=250\]n = 7 {2*7+1,2*7+3} -> {15,17}